# Stability of quasicrystalline ultracold fermions to dipolar interactions

### v1

@@ Abstract
Conclusions — Overall, we observed that strong DDI in quasiperiodic potentials not only inhibit spreading in the density, but also in the two-body correlations. This leads to a stabilization of localized states, which favors the intermediate phase for shallower potentials, and localized phases at deeper potentials. In particular, we observed stable, DDI-induced localization even in the absence of a detuning potential. When the detuning potential is gradually turned on, the corresponding localized states undergo resonance-like oscillations at the boundaries before returning to a fully localized behavior, indicating a rich interplay between DDI and quasiperiodicity.
Repulsive composite objects have been shown to be stable with respect to on-site interactions in a structured environment such as a simple optical lattice Winkler et al. (2006); Strohmaier et al. (2010); Schneider et al. (2012); Ronzheimer et al. (2013); Zaletel et al. (2015). Dynamically bound dimers are also been observed in extended Bose-Hubbard models Li et al. (2020, 2021). Our study propels these notions to a full many-body horizon by combining quasiperiodicity and dipolar interactions in a realistic quantum simulator setting Hughes et al. (2023); Bilinskaya et al. (2024), and should provide a guideline to explore exotic localization phenomena that combine long-range interactions and correlated disorder. Our scheme, inspired by experiments in 1D geometries, could be readily extended to higher-dimensional setups where a richer landscape of quasicrystalline phases with different rotational symmetries can be engineered Viebahn et al. (2019); Sbroscia et al. (2020); Gottlob and Schneider (2023).
@@ I System parameters
[…]
@@ I.1 Lengths
In MCTDH-X simulations, we choose to set the unit of length $\bar{L}\equiv\frac{\lambda_{p}}{2}=266.1$ nm, which makes the maxima of the primary lattice appear at integer values in dimensionless units. We run simulations with 2’048 gridpoints in an interval $x\in[-32\bar{L},32\bar{L}]\approx[-8.5152\mu\mathrm{m},8.5152\mu\mathrm{m}]$, giving a resolution of around 8.3156 nm. We employ hard-wall boundary conditions at the end of the probed spatial interval.
@@ I.2 Energies
The unit of energy is defined in terms of the recoil energy of the primary lattice, i.e. $E_{r}\equiv\frac{\hbar^{2}k_{p}^{2}}{2m}\approx 1.1679\times 10^{-29}$ J with $m\approx 39.9640$ Da the mass of ${}^{40}$ atoms. More specifically, we define the unit of energy as $\bar{E}\equiv\frac{\hbar^{2}}{m}\frac{1}{\bar{L}^{2}}=\frac{2E_{r}}{\pi^{2}}\approx 0.237\times 10^{-19}$ J. In typical experiments with quasiperiodic optical lattices, the depths are varied in regimes of up to around 8 recoil energies for the primary lattice, and around 1 recoil energy for the detuning lattice. In our simulations, we probe similar regimes: $V_{p}\in[15\bar{E},60\bar{E}]\approx[3.0E_{r},12.2E_{r}]$, $V_{d}\in[0\bar{E},6\bar{E}]=[0E_{r},1.2E_{r}]$. When we turn on dipolar interactions, we probe the regimes with strength $W\in[0\bar{E},10.0\bar{E}]\approx[0E_{r},2E_{r}]$, which is in accordance to the values that can be achieved in near-term ultracold dipolar quantum simulators.

### v2

@@ Abstract
Discussion — There are several mechanisms at play that explain the overall shape of the phase diagram at strong DDI. The localized state appearing at large $V_{p}$ (lattice limit) in the clean limit $V_{d}\to 0$ is a DSBS stabilized by DDI. The $1/r^{3}$ tails of the DDI cause an interference that locks particles within a certain distance, as demonstrated in similar bosonic systems Barbiero et al. (2015); Li et al. (2020, 2021a, 2021b); Korbmacher et al. (2023), and is likely connected with Hilbert space fragmentation Khemani et al. (2020); Herviou et al. (2021); Cheng (2023). Away from the clean limit ($V_{d}\neq 0$, but $V_{d}\ll V_{p}$), correlated disorder can be regarded as a perturbation on top of the DSBS. In this setting, it is known that the DDI give rise to effective correlated hopping at a distance, i.e. between the edges of the localized clusters Levitov (1989, 1990); Burin and Kagan (1994); Burin (2015a, b, 2017); Nosov et al. (2019), which explains the resonant-like behavior at the edges at their progressive delocalization. Increasing the disorder further eventually pushes the system again into localization. In this limit, the system approaches an Aubry-André model, whose localized phase is known to be stable against DDI Gopalakrishnan (2017); Nag and Garg (2019); Deng et al. (2019).
On the other hand, as $V_{p}$ is decreased, the lattice becomes shallower, increasing the effective particle hopping. In the clean limit ($V_{d}\to 0$), this progressively reduces the critical distance under which particles can be confined by the DDI Barbiero et al. (2015) and full localization is lost. With correlated disorder ($V_{d}\gg 0$), the loss of localization is analogous to what happens in the noninteracting case [cf. Fig. 1(a)]. However, the repulsive nature of the DDI – unhindered by the DSBS resonances in shallow lattices – provides more mobility, resulting in an enlarged intermediate phase.
Repulsive composite objects have been shown to be stable with respect to on-site interactions in a structured environment such as a simple optical lattice Winkler et al. (2006); Strohmaier et al. (2010); Schneider et al. (2012); Ronzheimer et al. (2013); Zaletel et al. (2015). Our study propels these notions to a full many-body horizon by combining quasiperiodicity and dipolar interactions in a realistic quantum simulator setting Hughes et al. (2023); Bilinskaya et al. (2024). Our scheme, inspired by experiments in 1D geometries, could be readily extended to higher-dimensional setups where a richer landscape of quasicrystalline phases with different rotational symmetries can be engineered Viebahn et al. (2019); Sbroscia et al. (2020); Gottlob and Schneider (2023). Moreover, clean system localization can appear when other mechanisms induce effective disorder, such as an external linear field Schulz et al. (2019); Yao and Zakrzewski (2020); Taylor et al. (2020); Doggen et al. (2021); Guo et al. (2021); Yao et al. (2021), slow interaction dynamics Yao et al. (2016), or both van Nieuwenburg et al. (2019). Given that the system presented in our work interpolates between the clean and the quasiperiodic case, it should help shed light on commonalities across a multitude of exotic localization phenomena that combine long-range interactions and correlated disorder.
@@ I System parameters
where $k_{i}=\frac{2\pi}{\lambda_{i}}$. We choose the wavelengths to be compatible with real experimental realizations in ultracold atomic labs, i.e. $\lambda_{p}\approx 532.2$ nm and $\lambda_{d}\approx 738.2$ nm. This gives wave vectors $k_{p}\approx 1.1806\times 10^{7}$ m-1, $k_{d}\approx 8.5115\times 10^{6}$ m-1. For the figures in the main text, we chose $\phi=4.0$. However, we performed calculations with various random value of $\phi$ and did not notice any qualitative difference in the dynamics.
@@ I.1 Lengths
In MCTDH-X simulations, we choose to set the unit of length $\bar{L}\equiv\frac{\lambda_{p}}{2}=266.1$ nm, which makes the maxima of the primary lattice appear at integer values in dimensionless units. We run simulations with 2’048 gridpoints in an interval $x\in[-32\bar{L},32\bar{L}]\approx[-8.5152\mu\mathrm{m},8.5152\mu\mathrm{m}]$, giving a resolution of around 8.3156 nm. We employ hard-wall boundary conditions at the end of the probe spatial interval.
@@ I.2 Energies
The unit of energy is defined in terms of the recoil energy of the primary lattice, i.e. $E_{r}\equiv\frac{\hbar^{2}k_{p}^{2}}{2m}\approx 1.1679\times 10^{-29}$ J with $m\approx 39.9640$ Da the mass of 40K atoms. More specifically, we define the unit of energy as $\bar{E}\equiv\frac{\hbar^{2}}{m}\frac{1}{\bar{L}^{2}}=\frac{2E_{r}}{\pi^{2}}\approx 0.237\times 10^{-19}$ J. In typical experiments with quasiperiodic optical lattices, the depths are varied in regimes of up to around 8 recoil energies for the primary lattice, and around 1 recoil energy for the detuning lattice. In our simulations, we probe similar regimes: $V_{p}\in[15\bar{E},60\bar{E}]\approx[3.0E_{r},12.2E_{r}]$, $V_{d}\in[0\bar{E},6\bar{E}]=[0E_{r},1.2E_{r}]$. When we turn on dipolar interactions, we probe the regimes with strength $W\in[0\bar{E},10.0\bar{E}]\approx[0E_{r},2E_{r}]$, which is in accordance to the values that can be achieved in near-term ultracold dipolar quantum simulators.